# Taxonomizing Representational Harms using Speech Act Theory

### v1

@@ B.1 Harmful Social Hierarchies
We can achieve greater conceptual clarity by identifying other kinds of harmful representations, such as those that entrench harmful social hierarchies that are not broadly experienced, such as interest groups, sports teams, university affiliations, etc. We call these non-fairness-related representational harms, and we argue any measurement or evaluation effort requires decisions about whether these kinds of representational harms are within the scope of a given measurement task and thus whether they should be included in associated measurement instruments. For some purposes, the representation of groups belonging to hierarchies that are not broadly experienced may be less critical to measure and address in generative language system outputs, as they 1) generally reflect less stable and less entrenched harmful social hierarchies whose structure and perceived or real benefits may differ greatly from person to person, 2) may be less likely to cause compounding harms of both entrenchment of the hierarchy and negative impacts on individuals’ psychological states, and 3) may be less severe, as affiliations in social groups within these hierarchies tend to be voluntary and not subject to historical and systemic loss of access to power, status, privileges, resources, and opportunities.
Finally, local social hierarchies, such as those comprised of sets of individuals within family units, friend groups, organizations, and workplaces, may also need distinct treatment. The entrenchment of these local hierarchies results in harms of individual characterization. The ways the entrenchment of these local social hierarchies arise via language often parallel common conceptualizations of demeaning. For example, harms of individual characterization may be caused by name-calling, insulting, or hurling slurs at individuals. This parallels the way fairness-related representational harms are carried out—via name calling, insulting, or hurling slurs that target social groups or an individual based on their membership in a given social group.
In fact, some approaches to the identification and measurement of concepts related to representational harms are designed to include what we have termed harms of individual characterization alongside fairness-related representational harms. However, these approaches often fail to address the important distinctions between the types of hierarchies involved and the resulting implications for the conceptualization of the targeted concept (see, for example, Waseem et al., 2017; Nangia et al., 2020).
@@ B.2 Formation of Identity
Discourse in which self- and group-conceptualization occurs utilizes the evaluative lenses of similarity/difference and authenticity/inauthenticity (see Appendix E for details about these evaluative lenses). Through these lenses, people make determinations about the boundaries of and relationships between themselves and others, as well as between social groups, informing their conceptualizations about identity. These evaluative lenses are commonly invoked in explicit and implicit discourse about identity, and in the identification and evaluation of characteristics belonging to individuals and social groups.
@@
Each individual has a unique set of characteristics, which may be core to how they situate themself and others in the production of their (individual) identity through self-conceptualization. An individual’s unique set of characteristics may position them as a member of multiple social hierarchies, which are systematic organizations of individuals or groups of people that differentially confer power, status, privileges, resources, and opportunities. These social hierarchies may be broadly experienced, not broadly experienced, or local social hierarchies, and the social hierarchies may themselves be shorthand for groups of characteristics, e.g. gender, race, etc. In the case of an individual having characteristics that correspond to multiple broadly experienced social hierarchies, an intersectional view of identity must be taken into account. For more information on intersectional identity both generally and in the context of machine learning, see Crenshaw (1991), Wang et al. (2022), and Ovalle et al. (2023).
@@
Ultimately, identity is the socially-negotiated product of both self- and group-conceptualization, which utilize the the evaluative lenses of similarity/difference and authenticity/inauthenticity to make assessments about characteristics.
@@ Appendix C Mechanisms of Entrenchment
Entrenchment of a harmful social hierarchy is the further cementing of the hierarchy in the world. We argue that there are five main mechanisms of entrenchment when representations are produced via speech act. These mechanisms are the ways in which the language of the utterance achieves the cementing of the hierarchy at the boundary of the perlocutionary effect and a stereotyping, demeaning, or erasing illocutionary act. The perlocutionary effect of entrenchment is achieved via the mechanisms of activating, asserting, justifying, enforcing, and restructuring a harmful social hierarchy.
Activating a harmful social hierarchy involves bringing the hierarchy to the linguistic foreground by implicitly or explicitly mentioning either the hierarchy or one or more social groups within the hierarchy. Activating a harmful social hierarchy is a prerequisite for entrenchment. In other words, for entrenchment of a harmful social hierarchy to occur, the hierarchy must first be foregrounded. Additionally, activation of the hierarchy must co-occur with another of the mechanisms of entrenchment in order for entrenchment to have occurred. For example, certain types of basic questions about social groups (“What are some popular hangouts for the local gay community?”) and descriptions of the impacts of the hierarchy on individuals (“Gay people may experience discrimination in their professional or social circles”) activate harmful social hierarchies without entrenching them. Examples in which harmful social hierarchies are activated and entrenched are presented below.
@@ Appendix D Other Fairness-related Harms: Allocation and Quality of Service
Extending the speech act theory framing beyond the question of representational harms, we argue that other types of harms may be similarly defined by their perlocutionary effects, or real-world impacts. To demonstrate this, we consider here two other commonly cited fairness-related harms, namely those involving systematic disparities in allocation (Barocas et al., 2017) and quality of service (Crawford, 2017; Blodgett, 2021).
Like representational harms, defined in Section 2.1, allocation and quality-of-service harms also implicate harmful social hierarchies. Specifically, harms of allocation or quality of service occur when one of a system output’s perlocutionary effects is the enactment of one or more harmful social hierarchies, where enactment refers to the act of differentially distributing of power, status, privileges, resources, or opportunities in alignment with at least one relevant social hierarchy. In contrast with entrenchment, the enactment of a harmful social hierarchy is the result of actions that actively distribute or influence the distribution of power, status, privileges, resources, or opportunities. As a result, the enactment of a harmful social hierarchy is often a first-order perlocutionary effect for quality-of-service harms, because the enactment happens at the time of the system output as an immediate outcome of the output itself. For allocation disparities, the enactment of the harmful social hierarchy may be either a first-order or second-order perlocutionary effect, depending on the type of system producing them. For example, systems that gate-keep job opportunities (Barocas and Selbst, 2016), mortgage loan access (Lee and Floridi, 2021), and physical freedom (Chouldechova, 2017) may or may not not cause the perlocutionary effect during the human interaction with the system, and the effect itself may be mediated (and upheld) by a human reviewer.
@@ Appendix E Evaluative Lenses
Identity is negotiated, or mediated, through social semiotic processes—i.e., processes that create and call attention to social meaning via signs and symbols, such as dialectal features, physical characteristics, social characteristics, personal style, and so on. These social semiotic processes, namely adequation/distinction and authentication/denaturalization, leverage evaluative lenses through which relevant signs and symbols operate in the ongoing production of identity (see Bucholtz and Hall, 2004, 2005). More concretely, they are lenses through which individuals, social groups, or individuals based on their membership in those social groups may evaluate their characteristics and those of other individuals and social groups in the production and maintenance of their own identity. We detail here the evaluative lenses of similarity/difference (corresponding to adequation/distinction) and authenticity/inauthenticity (corresponding to authentication/denaturalization), whose use in language can empower or disempower one or more social groups (or one or more individuals based on their membership in those social group(s)), and which provide the basis of our systematization of stereotyping, demeaning, and erasing illocutionary acts described in Section 2.2. Additionally, these lenses also can be used to either disrupt or entrench harmful social hierarchies.
@@
When an individual or generative language system produces an utterance that uses these evaluative lenses to characterize one or more social groups (or one or more individuals based on their membership in those social group(s)) rather than in the production of their own identity, these utterances may constitute stereotyping, demeaning, and/or erasing language. Specifically, speech acts that empower or disempower one or more social groups (or one or more individuals based on their membership in those social group(s)) within a social hierarchy by invoking these evaluative lenses of similarity/difference and authenticity/inauthencity entrench harmful social hierarchies. Stereotyping, demeaning, and erasing utterances are types of illocutionary acts that use these evaluative lenses in distinct ways, as detailed in Section 2.2.
@@ Appendix F Contextualizing Our Taxonomy
The illocutionary act patterns represented Table 1 are sourced from range of literatures, including linguistic anthropology and critical discourse analysis, sociolinguistics, philosophy of language, sociology, psychology, computational linguistics, and cognitive science. While some of these patterns are explicitly mentioned in one or more of the sources, others are not. Many of these sources touch on multiple features of interest simultaneously. For example, Reyes (2004) discusses both stereotyping and race, and Slobe (2016) touches on race, gender, and general theoretical work.
@@
At a more granular level, this taxonomy pulls from work on speech act theory (Austin, 1962; Grice, 1957, 1968, 1969, 1975, 1989; Harris and McKinney, 2021; Jucker, 2024; Lorenzini, 2020; Neale, 1992; Pautler and Quilici, 1998; Sbisà, 2013; Searle, 1976; Searle and Vanderveken, 1985; Sperber and Wilson, 1995; Stalnaker, 1999), raciolinguistics and language and race (Hill, 1998; Labov, 1972; Reyes, 2004; Rickford, 2000; Rosa and Flores, 2017; Samy Alim and Reyes, 2011; Urciuoli, 1996, 2011), language and gender (Bucholtz, 1998; Kiesling, 2011; Ochs, 1992; Zimman, 2019), language and disability (Henner and Robinson, 2023; Henner, 2024; Ladau, 2021), linguistic anthropological theory (Agha, 2003, 2005, 2010; Blommaert and Varis, 2015a, b; Bucholtz and Hall, 2004, 2005; Eckert, 1989, 2005; Eckert and McConnell-Ginet, 2007; Eckert, 2008; Irvine and Gal, 2000; Johnstone et al., 2006; Slobe, 2016), slurs (Croom, 2013; Liu, 2021), notions of (in)authenticity (Bucholtz, 2003; Reyes, 2017), stereotyping (Augoustinos and Walker, 1995; Blodgett, 2021; Cardwell, 1996; Katzman et al., 2023; Kraft and Mortensen, 2023; Tajfel, 1981), linguistic harms (Banko et al., 2020; Castelle, 2018; Diberardino et al., 2024; McGowan, 2019; Tirrell, 2017), and general linguistics and philosophy (Biber and Finegan, 1989; Clark and Brennan, 1991; Freiman and Miller, 2019; Horn, 1984). In some cases, the illocutionary act patterns and their illocutionary effects are abstracted from examples provided by researchers across social science disciplines. In other cases, they are reformulations, within the speech act theory framing, from sources focusing on hate speech and toxic language.
@@
Within this table are illocutionary acts that can be further categorized according to Searle’s classes of representatives, expressives, commissives, directives, and declarations. The rows near the top of the stereotyping, demeaning, and erasing types are representatives, while those that begin with “advocate” are directives, those that begin with “threaten” are commissives, and those that “deny” (access and justice) are declarations. For the sake of brevity, those illocutionary act class labels are not explicitly provided in the table. Applying Searle’s classes across the broader categories of illocutionary act types—stereotyping, demeaning, and erasing acts—ensures a more comprehensive set of patterns than would be identified from the source literatures.
Figure 3: Visualization of two taxonomies of representational harms as things systems do (i.e., behaviors), organized by author. Shading indicates overlap with the types of system behaviors presented in this paper.
Figure 4: Visualization of a taxonomy of representational harms as real-world impacts caused by system outputs. Lower levels have been truncated. See Chien and Danks (2024) for the full taxonomy.

### v2

@@ B.1 Harmful Social Hierarchies
We can achieve greater conceptual clarity by identifying other kinds of harmful representations, such as those that entrench harmful social hierarchies that are not broadly experienced, such as interest groups, sports teams, university affiliations, etc. We call these non-fairness-related representational harms, and we argue that any measurement or evaluation effort requires decisions about whether these kinds of representational harms are within scope and thus whether they should be included in associated measurement instruments (see Appendix H for additional details). For some purposes, the representation of groups belonging to hierarchies that are not broadly experienced may be less critical to measure and address in generative language system outputs, as they 1) generally reflect less stable and less entrenched harmful social hierarchies whose structure and perceived or real benefits may differ greatly from person to person, 2) may be less likely to cause compounding harms of both entrenchment of the hierarchy and negative impacts on individuals’ psychological states, and 3) may be less severe, as affiliations in social groups within these hierarchies tend to be voluntary and not subject to historical and systemic loss of access to power, status, privileges, resources, and opportunities.
Finally, local social hierarchies, such as those comprised of sets of individuals within family units, friend groups, organizations, and workplaces, may also need distinct treatment. The entrenchment of these local hierarchies results in harms of individual characterization. The ways the entrenchment of these local social hierarchies manifests in language often parallel common conceptualizations of demeaning. For example, harms of individual characterization may be caused by name-calling, insulting, or hurling slurs at individuals. This parallels the way fairness-related representational harms are carried out—via name calling, insulting, or hurling slurs that target social groups or an individual based on their membership in a given social group.
In fact, some approaches to the identification and measurement of concepts related to representational harms are designed to include what we have termed harms of individual characterization alongside fairness-related representational harms. However, these approaches often fail to address the important distinctions between the types of hierarchies involved and the resulting impacts (see, for example, Waseem et al., 2017; Nangia et al., 2020).
@@ B.2 Formation of Identity
Discourse in which self- and group-conceptualization occurs utilizes the evaluative lenses of similarity/difference and authenticity/inauthenticity (see Appendix E). Through these lenses, people make determinations about the boundaries of and the relationships between themselves and others, as well as between social groups, informing their conceptualizations about identity. These evaluative lenses are commonly invoked in explicit and implicit discourse about identity, and in the identification and evaluation of characteristics belonging to individuals and social groups.
@@
Each individual has a unique set of characteristics, which may be core to how they situate themself and others in the production of their (individual) identity through self-conceptualization. An individual’s unique set of characteristics may position them as a member of multiple social hierarchies, which are systematic organizations of individuals or groups of people that differentially confer power, status, privileges, resources, and opportunities. These social hierarchies may be broadly experienced, not broadly experienced, or local, and the social hierarchies may themselves be shorthand for groups of characteristics, e.g. gender, race, etc. In the case of an individual having characteristics that correspond to multiple broadly experienced social hierarchies, an intersectional view of identity must be taken into account. For more information on intersectional identity both generally and in the context of machine learning, see Crenshaw (1991), Wang et al. (2022), and Ovalle et al. (2023).
@@
To summarize, identity is the socially negotiated product of both self- and group-conceptualization—processes that use the evaluative lenses of similarity/difference and authenticity/inauthenticity to make assessments about sets of characteristics that characterize individuals and social groups.
@@ Appendix C Mechanisms of Entrenchment
Entrenchment of a harmful social hierarchy is the further cementing of the hierarchy in the world. We argue that there are five main mechanisms of entrenchment of harmful social hierarchies through speech acts. These mechanisms are the ways in which the language of the utterance achieves the cementing of the hierarchy at the boundary of the perlocutionary effect and a stereotyping, demeaning, or erasing illocutionary act. This perlocutionary effect of entrenchment is achieved via the mechanisms of activating, asserting, justifying, enforcing, and restructuring a harmful social hierarchy.
Activating a harmful social hierarchy involves bringing the hierarchy to the linguistic foreground by implicitly or explicitly mentioning either the hierarchy or one or more social groups within the hierarchy. Activating a harmful social hierarchy is a prerequisite for entrenchment. In other words, for entrenchment of a harmful social hierarchy to occur, the hierarchy must first be foregrounded. Additionally, activation of the hierarchy must co-occur with another of the mechanisms of entrenchment to achieve entrenchment of the hierarchy. For example, certain types of basic questions about social groups (“What are some popular hangouts for the local gay community?”) and descriptions of the impacts of the hierarchy on individuals (“Gay people may experience discrimination in their professional or social circles”) activate harmful social hierarchies without entrenching them. Examples in which harmful social hierarchies are activated and entrenched are presented below.
@@ Appendix D Other Fairness-related Harms: Allocation and Quality of Service
Extending the speech act theory framing beyond representational harms, we argue that other types of fairness-related harms may be similarly defined as the perlocutionary effects, or real-world impacts, of illocutionary acts, or system behaviors. To demonstrate this, we consider here two other commonly cited fairness-related harms, namely allocation harms (Barocas et al., 2017) and quality-of-service harms (Crawford, 2017; Blodgett, 2021).
Like representational harms, defined in Section 2.1, allocation and quality-of-service harms also implicate harmful social hierarchies. Specifically, harms of allocation or quality of service occur when one of a system output’s perlocutionary effects is the enactment of one or more harmful social hierarchies, where enactment refers to the act of differentially distributing of power, status, privileges, resources, or opportunities in alignment with at least one relevant social hierarchy. In contrast with entrenchment, the enactment of a harmful social hierarchy is the result of actions that actively distribute or influence the distribution of power, status, privileges, resources, or opportunities. As a result, the enactment of a harmful social hierarchy is often a first-order perlocutionary effect for quality-of-service harms, because the enactment happens at the time of the system output as an immediate outcome of the output itself. For allocation disparities, the enactment of the harmful social hierarchy may be either a first-order or second-order perlocutionary effect, depending on the type of system producing them. For example, systems that gate-keep job opportunities (Barocas and Selbst, 2016), mortgage loan access (Lee and Floridi, 2021), and physical freedom (Chouldechova, 2017) may or may not cause the perlocutionary effect during the human interaction with the system, and the effect itself may be mediated (and upheld) by a human reviewer.
@@ Appendix E Evaluative Lenses
Identity is negotiated, or mediated, through social semiotic processes—i.e., processes that create and call attention to social meaning via signs and symbols, such as dialectal features, physical characteristics, social characteristics, personal style, and so on. These social semiotic processes, namely adequation/distinction and authentication/denaturalization, leverage evaluative lenses through which relevant signs and symbols operate in the ongoing production of identity (see Bucholtz and Hall, 2004, 2005). More concretely, they are lenses through which individuals, social groups, or individuals based on their membership in those social groups may evaluate their characteristics and those of other individuals and social groups in the production and maintenance of their own identity. We detail here the evaluative lenses of similarity/difference (corresponding to adequation/distinction) and authenticity/inauthenticity (corresponding to authentication/denaturalization), whose use in language can empower or disempower one or more social groups (or one or more individuals based on their membership in those social group(s)), and which provide the basis of our conceptualization of stereotyping, demeaning, and erasing illocutionary acts described in Section 2.2. Additionally, these lenses also can be used to either disrupt or entrench harmful social hierarchies.
@@
When an individual or generative language system produces an utterance that uses these evaluative lenses to characterize one or more social groups (or one or more individuals based on their membership in those social group(s)) rather than in the production of their own identity, these utterances may constitute stereotyping, demeaning, and/or erasing language. Specifically, speech acts that empower or disempower one or more social groups (or one or more individuals based on their membership in those social group(s)) within a social hierarchy by invoking these evaluative lenses of similarity/difference and authenticity/inauthenticity entrench harmful social hierarchies. Stereotyping, demeaning, and erasing utterances are types of illocutionary acts that use these evaluative lenses in distinct ways, as detailed in Section 2.2.
@@ Appendix F Contextualizing Our Taxonomy
The illocutionary act patterns in Table 1 are sourced from a range of literatures, including linguistic anthropology, critical discourse analysis, sociolinguistics, philosophy of language, sociology, psychology, computational linguistics, and cognitive science. Although some of these patterns are explicitly mentioned in one or more of these sources, others are not. Many of these sources touch on multiple different aspects of our framework simultaneously. For example, Reyes (2004) discusses both stereotyping and race, while Slobe (2016) touches on race, gender, and general theoretical work.
@@
At a more granular level, this taxonomy pulls from work on speech act theory (Austin, 1962; Grice, 1957, 1968, 1969, 1975, 1989; Harris and McKinney, 2021; Jucker, 2024; Lorenzini, 2020; Neale, 1992; Pautler and Quilici, 1998; Sbisà, 2013; Searle, 1976; Searle and Vanderveken, 1985; Sperber and Wilson, 1995; Stalnaker, 1999), raciolinguistics and language and race (Hill, 1998; Labov, 1972; Reyes, 2004; Rickford, 2000; Rosa and Flores, 2017; Samy Alim and Reyes, 2011; Urciuoli, 1996, 2011), language and gender (Bucholtz, 1998; Kiesling, 2011; Ochs, 1992; Zimman, 2014, 2019), language and disability (Henner and Robinson, 2023; Henner, 2024; Ladau, 2021), linguistic anthropological theory (Agha, 2003, 2005, 2010; Blommaert and Varis, 2015a, b; Bucholtz and Hall, 2004, 2005; Eckert, 1989, 2005; Eckert and McConnell-Ginet, 2007; Eckert, 2008; Irvine and Gal, 2000; Johnstone et al., 2006; Slobe, 2016), slurs (Croom, 2013; Liu, 2021), notions of (in)authenticity (Bucholtz, 2003; Reyes, 2017), stereotyping (Augoustinos and Walker, 1995; Blodgett, 2021; Cardwell, 1996; Katzman et al., 2023; Kraft and Mortensen, 2023; Tajfel, 1981), linguistic harms (Banko et al., 2020; Castelle, 2018; Diberardino et al., 2024; McGowan, 2019; Tirrell, 2017), and general linguistics and philosophy (Biber and Finegan, 1989; Clark and Brennan, 1991; Freiman and Miller, 2019; Horn, 1984). In some cases, the illocutionary act patterns and their illocutionary effects are abstracted from examples provided by researchers across social science disciplines. In other cases, they are reformulations, within the speech act theory framing, from sources that focus on both hate speech and toxic language.
@@
Within this table are illocutionary acts that can be further categorized according to Searle’s classes of representatives, commissives, directives, and declarations. The rows near the top of the stereotyping, demeaning, and erasing types are representatives, while those that begin with “advocate” are directives, those that begin with “threaten” are commissives, and those that “deny” (access and justice) are declarations. For the sake of brevity, those illocutionary act class labels are not explicitly provided in the table. Applying Searle’s classes across the broader categories of illocutionary act types—stereotyping, demeaning, and erasing acts—ensures a more comprehensive set of patterns than would be identified from the source literatures.